OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = 9 –3x + 2y = 9

OpenStudy (anonymous):

please help me do this thanks

OpenStudy (anonymous):

( -6,9) i think

OpenStudy (anonymous):

but do you know how?

OpenStudy (anonymous):

no thats why im here to ask for help

OpenStudy (anonymous):

A. (–4, 3) B. (0, 6) C. (3, 4) D. (4, 3)

OpenStudy (anonymous):

can someone show me how to do this please

OpenStudy (anonymous):

Here let me show you the steps

OpenStudy (anonymous):

3x – 4y = 9 –3x + 2y = 9 First, just cross out the x's since they are already opposites. -2y = 18 y = -9 3x - 4(-9) = 9 3x + 36 = 9 3x = -27 x = -9

OpenStudy (anonymous):

so the answer would be c??

OpenStudy (anonymous):

@Brybro ???

OpenStudy (anonymous):

What is C? Your list of answers up top are all wrong. Even if you plug them into the equations, they don't work. Only (-9, -9) works.

OpenStudy (anonymous):

hmmm